OpenStudy (anonymous):

The first term of an arithmetic sequence is -3 and the fifteenth term is 53. What is the common difference of the sequence? a.) 14/13 b.) 25/7 c.)4

OpenStudy (campbell_st):

well a term in an arithmetic sequence can be found using the formula \[a_{n} = a + (n -1) \times d\] you know n = 15 and a = -3 and a3 = 53 substitute and then solve for d